OpenStudy (anonymous):

Trig Identities: Prove. (sin^3x + cos^3x)/sinx +cosx= 1 -sinxcosx

OpenStudy (anonymous):

is it like this? \[ \frac{\sin^3x+\cos^3x}{\sin x}+\cos x \]

OpenStudy (anonymous):

no the sinx + cosx are both a the bottom

OpenStudy (anonymous):

\[ \frac{\sin^3x+\cos^3x}{\sin x+\cos x}\\ \quad=\frac{\sin x(1-\cos^2x)+\cos x(1-\sin^2x)}{\sin x+\cos x}\\ \quad=\frac{\sin x-\cos^2x\sin x+\cos x-\sin^2x\cos x}{\sin x+\cos x}\\ \quad=\frac{(\sin x+\cos x)-\sin x\cos x(\sin x+\cos x)}{\sin x+\cos x}\\ \quad=\frac{\cancel{(\sin x+\cos x)}(1-\sin x\cos x)}{\cancel{\sin x+\cos x}} \]

OpenStudy (anonymous):

okay i see... so in the last step when you cross them out that is the one? @electrokid

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is the one the entire thing should equal to right? so, QED.

OpenStudy (anonymous):

Thank You!